OpenStudy (anonymous):

5x + y = 6 2y = x +1

OpenStudy (anonymous):

x=1 y=1

OpenStudy (anonymous):

lOok : we can have : 2y=x+1=2y-x=1 Right ?:)

OpenStudy (anonymous):

x + y = 3 2x - y = 3

OpenStudy (anonymous):

Is it another question ????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Did u get last ????

OpenStudy (anonymous):

yes i did E.ali

OpenStudy (anonymous):

in this just + Xs together , numbers together and Ys togrther ...

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Means : x + y = 3 2x - y = 3 x+2x y-y 3+3

OpenStudy (anonymous):

so it will equal (1,1)

OpenStudy (anonymous):

No on ! Look : x+2x=3x y-y=0 3+3=6 So we have : 3x=6 Got it ???:)

OpenStudy (anonymous):

got it (2,1)

OpenStudy (anonymous):

_1 or 1 ?? Caculate y again please ..:)

OpenStudy (anonymous):

no im tell u not ask the right answer was (2,1)

OpenStudy (anonymous):

Yes ! You re realy right !;)